OpenStudy (anonymous):

Let f(x)=x^3-7x^2+25x-39 and let g be the inverse function of f. What is the value of g'(0)?

OpenStudy (nottim):

I guess you should do the inverse of the equation first?

OpenStudy (anonymous):

can someone pleaseeee help me????????

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

I think I have the same question...hold on

OpenStudy (nottim):

you didnt leanr inverse?

OpenStudy (anonymous):

Inverse of the equation would take too much work. Consider the following: let f and g be as stated. Then since they are inverses, clearly f(g(x))=x. Taking the derivative of each side, you get \[f'(g(x))g'(x)=1\] Thus, \[g'(x)=\frac{1}{f'(g(x))}\] Thus, \[g'(0)=\frac{1}{f'(g(0))}\] Since g(0) is just the point x such that f(x)=0. A little algebra shows the only real answer to this is x=3. Since \[f'(x)=3x^2-14x+25\] you clearly have that \[f'(3)=27-42+25=10\] Hence, you have the final answer that \[g'(0)=\frac{1}{10}\]

OpenStudy (anonymous):

not completely

OpenStudy (anonymous):

Thank you imperialist